OpenStudy (anonymous):

If a=5 and c=10 use the pythagorean theorem to solve for b.

OpenStudy (anonymous):

\[5\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 5^2 + b^2 = 10^2 25 +b^2 = 100 b^2 = 100-25 b^2 = 75 b = sqrt(75) b = sqrt(25*3) b = sqrt(25)*sqrt(3) b = 5*sqrt(3)

OpenStudy (anonymous):

First off. Which value is the length of the hypotenuse?

OpenStudy (lgbasallote):

@Grazes it would be most appreciated if you would not only post answers only :) the asker would appreciate it if you would give her a solution yes? Furthermore, it is against the site's Code of Conduct to do so. ( http://openstudy.com/code-of-conduct). Don't worry you're not yet in trouble ^_^ I am just telling you to avoid future mistakes, okay? Thank you :)

OpenStudy (anonymous):

I never read the code of conduct... They're usually really long.

OpenStudy (lgbasallote):

hahaha it's okay ^_^ all you need to know is don't be rude; be nice, and that this is an educational site so answers and questions must be educational..you'll get it as time goes by don't worry :DDD